Find the consumer's surplus at $P = 5$ for the following demand functions:

(a) P=25-2Q,

(b) P= 10/sqrt {Q} 


1
Expert's answer
2022-03-29T17:23:35-0400

a)

"P=25-2Q"

but "P=5"

"5=25-2Q"

"2Q=20"

"Q=10"

Consumer surplus "=\intop^Q_0D(Q)dQ-PQ"

"\intop^{10}_0(25-2Q)dQ-5(10)"

"=[25Q-Q^2]_0^{10}-50"

"=250-100-50"

"=100"


b)

"P=\frac{10}{\sqrt{Q}}"

"P=5"

"5=\frac{10}{\sqrt{Q}}"

"\sqrt{Q}=2"

"Q=4"

Consumer surplus "=\intop^4_0\frac{10}{\sqrt{Q}}dQ-20"

"=10[2\sqrt{Q}]^4_0-20"

"=20(\sqrt{4}-0)-20"

"=40-20"

"=20"